Controls Meeting 20-Mar-2014
- Minutes from previous meeting Controls Meeting 6-Mar-2014
- UPS requirements for Hall D (Dave, Yi, Hovanes)
- BCAL power supply interlocks (Dave)
- PLC archiving
- Radiator insertion stage test (Hovanes, Yi)
- Suggested modifications to BCAL and FDC voltage software (Elton, others)
- Any other business
9:30AM Thurs 20-Mar-2014 CC A110
- Dave collected the information from Jonathan regarding the power for the valves
- Yi would like to have the PXI system on the UPS as well (rated 300W)
- Hovanes showed the layout of the UPS requirements in Hall D (PXI was missing). Hovanes will send the requirements to Joe Beaufait on Friday.
BCAL power supply interlocks
- Dave showed slides explaining his work on the BCAL interlocks.
- Encountered strange problem that the 5V output modules on PoinIO did not provide enought current for the MPOD interlock input.
- Dave will order CompactLogix extension modules to avoid PintIO all together, and use CompactLogix modules.
- There was a request to change the button labels to adhere to Elton's convention.
- Hovanes says that Paul Letta found that accumulated 12GB of data in the PLC_ARCHIVE_ON_CHANGE database. This database was active about one month. Therefore we need to reduce the rate at which this DB gets filled by PLC.
- According to Dave, PLC only archived data for a short period of time, but it was not set up to be on-change.
- Hovanes will remove the data from the PLC_ARCHIVE_ON_CHANGE table, and Dave will work on on-change archiving setup.
Radiator stage test
- Hovanes said that the radiator insertion stage has been installed on the laser table in room 213 in FEL by Yi, Slava and himself, the same palce where the goniometer was tested. Scot and Tina made the cables to connect the stage to the XPS controller. Using a laser system the stage will be tested.
Suggested modification to voltage controls
- Elton said that he will e-mail a list of suggestion for improvements the BCAL voltage controls software (see Elton's future e-mail below).
Hi Nerses, I have been compiling feedback on the BCAL HV control 1. Would like to highlight the channel that has been selected. At the moment the channel number is displayed, but it would be nice to have a visual clue to point to the selected channel 2. The "show bias channels" window will not allow the voltage to be set above 70 V. The "set bias channels" window does not have this restriction. 3. The histogram display of the current is not operational 4. In the "set bias channels" window the measured temperature seems to display the measured voltage. 5. The "red" status on the overall display will remain ON even if some of the HV channels are on. The status will only turn green if all four HV channels from a given module are all on. For example, if one of the four HV channels from one module is off, that module will show "red". However, it is good to be able to look at the overall display and see if ANY channels are on. We may need to add another color, e.g. orange, that would show that some of the HV channels are on. * Finally, the issue of some channels staying at zero volts, even if the set voltage is 75 V. Yesterday I found this problem for BCAL module 13, which is a completely different HV module from previously identified module. I will continue to debug this; in the previous case, the channels were cleared when we unplugged the IPOD crate and rebooted the IOC. Thanks, Elton.
Dave will order CompactLogix extensions with 5V output modules for BCAL interlocks.
- Dave will modify the buttons on the BCAL interlocks to match Elton's convention.
Hovanes will send the UPS requirements to Joe Beaufait (after including Yi's PXI system requirements).
- Hovanes will remove all the data from PLC_ARCHIVE_ON_CHANGE.
Elton will send an e-mail with his suggestions for BCAL voltage controls improvements.